Vernon Corporation operates three investment centers. The following financial statements apply to the investment center named Bowman Division:

BOWMAN DIVISION
Income Statement
For the Year Ended December 31, 2018
Sales revenue $ 105,480
Cost of goods sold 59,475
Gross margin 46,005
Operating expenses
Selling expenses (2,750 )
Depreciation expense (4,045 )
Operating income 39,210
Nonoperating item
Loss of sale of land (3,800 )
Net income $ 35,410

BOWMAN DIVISION
Balance Sheet
As of December 31, 2018
Assets
Cash $ 12,532
Accounts receivable 40,436
Merchandise inventory 36,900
Equipment less accumulated depreciation 90,348
Nonoperating assets 10,300
Total assets $ 190,516
Liabilities
Accounts payable $ 9,557
Notes payable 70,000
Stockholders’ equity
Common stock 77,000
Retained earnings 33,959
Total liabilities and stockholder's equity $ 190,516

Required

  1. Calculate the ROI for Bowman.

  2. Vernon has a desired ROI of 13 percent. Headquarters has $92,000 of funds to assign to its investment centers. The manager of the Bowman Division has an opportunity to invest the funds at an ROI of 15 percent. The other two divisions have investment opportunities that yield only 14 percent. Calculate the new ROI for Bowman division, if the investment opportunity is adopted by Bowman.

  3. Based on the original data, calculate the original residual income. Also, calculate the new residual income based on information provided in Requirement d.
Answer #1
Part-1
1. ROI= Net Operatig Income / Net Operating Investment
39210/$180216= 21.76%
*190516-10300=$180216
Part-2
Adition Income on Investment = $92000*15%=$13800
New ROI= (Current Net Income + Additional Income earn on New Investment)/ Revised Net Operating Asset
(39210+13800)/ (180216+92000)= 19.47%
Part-3
Residual Income = Net Operating Income - (Average Operating Asset X Minimum Required return )
=$39210-( $180216X 13%)=$15782
Revised Residual Income = Revised Net Operating Income - (Revised Average Operating Asset X Minimum Required return )
=$53010-( $272216X 13%)=$17622